# A Systematic Literature Review of Spatio-Temporal Graph Neural Network Models for Time Series Forecasting and Classification

### v1

@@ Abstract
In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture dependencies among variables and across time points. The objective of the presented systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting. A database search was conducted, and over 150 journal papers were selected for a detailed examination of the current state-of-the-art in the field. This examination is intended to offer to the reader a comprehensive collection of proposed models, links to related source code, available datasets, benchmark models, and fitting results. All this information is hoped to assist researchers in future studies. To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of the results of current spatio-temporal GNN models in different domains. In addition, in its final part this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
@@ 1 Introduction
In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models for processing data that can be represented as graphs. GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks, image analysis, and natural language processing. The intuitive concept behind this approach is that nodes in a graph can be put in correspondence to objects or concepts, whereas edges can be assumed to represent their mutual relationships. Data for these quantities can then be processed at once by a GNN model. GNNs can be used for three different classes of problems: at graph level, at edge level, and at node level. In a graph-level problem, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges. For example, a molecule from a sample can be represented as a graph, and global properties of this molecule can be inferred from data coming from its entire structure in relation to the sample. In an edge-level problem, the goal is to predict the presence or absence of edges between pairs of nodes. As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions. In node-level problems, the goal is to predict the identity or role of each node within a graph, in order to solve either a classification or a regression task. Noticeably, regression tasks can be ”autoregressions”, that is, they can include time.
@@
This review examines then the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting. The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once. This results in a spatio-temporal GNN approach, where the spatial dimension is related to the multivariate framework, and the temporal dimension is related to the temporal nature of the data. Namely, spatio-temporal GNNs are a class of GNN models designed to handle data in both spatial and temporal dimensions at once. In this case, in a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is here to model the related dynamic interactions over time.
@@
Notwithstanding all this potential interest, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications. Most of the existing surveys are not SLRs, they are not recent, or only focus on a limited number of application domains. Individually, recent reviews focus for example on algorithmic features [cit_rev1], [cit_rev2], [cit_rev3], [survey_Lucheroni], GNN characterization [cit_rev8], specific fields [cit_rev4], and distributed training [cit_rev5], [cit_rev6]. Only a SLR on spatio-temporal GNNs was indeed recently published by Zeghina et al. [survey_spatiotemporal], but due to the restrictive search query, it covers only 52 publications.
@@
The intention of the SLR presented here is to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas. This SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields. Additionally, it collects results and benchmarks to assist researchers in their work by synthesizing the available literature on GNN applications. To the best of our knowledge, this is the first review to present comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature. It is thus hoped that this collected and distilled knowledge can become a valuable resource and reference for researchers.
@@
In this review two sets of questions will be posed. The answers to the first set want to provide a general overview of the publications in the field, while the answers to the second set focus on specific aspects of the proposed GNN models.
@@ 1. GQ1)
    Trend: What is the temporal trend of publications on GNN models? Is there a growing interest in GNNs?
@@ 3. GQ3)
    Journals: In which journals are the papers on GNNs published? Are these journals chosen because of the specific domain of application or not?
@@ 3. SQ3)
    Taxonomy: Among the various taxonomy classes, which are the most common? Are there some recurring mechanisms?
@@ 4. SQ4)
    Benchmark models: Which are the most common benchmark models? Are they classical ML models or also GNN models?
@@ 5. SQ5)
    Benchmark datasets: Which are the most common benchmark datasets?
@@ 6. SQ6)
    Modeling paradigms: What is the most common paradigm? Is it modeling complex interacting systems, or modeling a system with multiple interacting quantities? Specifically, do graphs typically condense the relationships among multiple entities, or do they describe different aspects of the same entity? In other words, is the graph structure homogeneous or heterogeneous?
@@ 7. SQ7)
The answers to both the general and the specific questions will be integrated into the following discussion in order to provide a comprehensive overview of the present state-of-the-art of spatio-temporal GNN models. A general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields. However, in case, for a more detailed analysis of a specific discipline or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before ([cit_rev1], [cit_rev2], [cit_rev3], [survey_Lucheroni], [cit_rev8], [cit_rev4], [cit_rev5], [cit_rev6]).
The paper is organized as follows. After this Introduction, Sec. 2 illustrates the methodology used to collect the papers. Sec. 3 provides a general overview of the selected papers. Sec. 4 introduces some fundamental definitions and notions at the basis of spatio-temporal GNN models, and discusses their taxonomy and the determination of the graph structure. Sec. 5 is the core of this review. It explores the diverse domains of application of the spatio-temporal GNN models proposed in the selected papers for the identified domains. A discussion of the findings and the answers to the research questions mentioned above are provided in Sec. 6. Finally, limits, challenges and future research directions are presented in Sec. 7. Sec. 8 concludes the review. Appendix  A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting).
@@ 2 Methodology for the systematic literature review
To conduct this SLR, four primary databases were consulted: Scopus, IEEE Xplore, Web of Science, and ACM. These databases were chosen for their comprehensive coverage of academic and scientific literature, ensuring a broad and inclusive search across multiple disciplines.
Prior to commencing the literature search, some exclusion criteria were defined. First, papers from the current year 2024 were excluded because it has not yet ended, hence the review only covers up to year 2023. Second, only journal articles were considered, in order to guarantee the inclusion of ”established” sources, whereas both conference papers and book chapters were excluded. The exclusion of conference papers is also made in order to keep the number of papers manageable. However, it is important to notice that relevant conference papers which provide useful benchmarks are anyway referenced throughout the text. Third, only articles published in English were included because of facility of fruition and comparison. Review articles are also not included, although they may be mentioned in the test.
An advanced search query submitted to the databases was designed to capture the breadth of the topic while maintaining specificity. The query used is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). The first two elements of query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
The selection process of the papers followed the multi-stage approach described below:
@@
    Data extraction: relevant data were extracted from the included studies, and analyzed in order to better discuss the approach and present the most common datasets and models for each identified field.
A total of 473 records were identified, 188 from Scopus, 31 from IEEE, 104 from Web of Science, and 150 from ACM. A scheme of this selection process is sketched in Fig. 1. First, 142 duplicates were removed, then the titles and abstracts of the remaining 331 records were manually evaluated to select the pertinent papers. Among these, 170 were discarded because they were considered outside the scope of this review, and 5 were discarded because they were not in English. Hence, 156 records were included in the review.
Figure 1: PRISMA flowchart to summarize the identification of the studies and the selection process, adapted from Ref. [PRISMA_statement].
Next, Sec. 3 provides an overview of the selected publications.
@@ 3 Overview of the publications
This section provides a comprehensive overview of the 156 selected journal articles. After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec. 5. In the following, some bibliographic information about the reviewed articles is presented.
Fig. 2 illustrates the number of publications over time across different groups. A cumulative positive trend can be observed over time, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020. The three most common groups that have seen the greatest increase in publications are ”Environment”, ”Generic”, and ”Mobility”. For more details, see also the group numerosity reported in Tab. 1.
@@
Field / Year 2020 2021 2022 2023 Total
Energy 1 6 6 13
Environment 2 4 21 27
Finance 3 6 9
Generic 1 4 19 24
Health 1 2 5 8
Mobility 2 4 15 18 39
Other 2 4 14 20
Predictive monitoring 2 3 11 16
Total 3 12 41 100 156
An inspection of the publication sources indicates that 95 different journals have published articles on GNN models for time series forecasting or classification. Tab. 2 lists the 15 journals which contain the highest number of published papers.
Table 2: Top 15 journals with the highest number of published paper.
Journal Number of papers
IEEE Access 7
ACM Transactions on Knowledge Discovery from Data 6
Applied Intelligence 6
IEEE Transactions on Instrumentation and Measurement 6
ACM Transactions on Intelligent Systems and Technology 5
Expert Systems with Applications 5
Sensors 4
Applied Energy 3
Applied Soft Computing 3
Energy 3
Engineering Applications of Artificial Intelligence 3
IEEE Transactions on Intelligent Transportation Systems 3
Information Sciences 3
International Journal of Data Science and Analytics 3
Knowledge-Based Systems 3
The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields. This suggests that researchers often select journals based on the specific domain of application, allowing them to reach targeted audiences and contribute within their respective fields. Such diversity also underlines the interdisciplinary nature of GNNs research, as it spans multiple domains.
@@
In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [VOSviewer]. Fig. 4 illustrates authors as nodes and collaborations as connections, revealing some research clusters. Two of these clusters are enlarged and shown in Fig. 4.
(a) Visualization of the complete collaboration network.
(b) Zoom in on two different clusters.
In the figure, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated. Two key observations can be made. First, the number of clusters suggests that there are many distinct research groups, indicating a high degree of fragmentation within the community. Second, the color of the nodes indicates that some groups were active only at the beginning and have since disappeared, while others have recently formed, as indicated by their yellow color. Some authors have various collaborations over time, as also shown in Fig. 4.
Fig. 3 shows a pie chart of the distribution of corresponding authors among different countries. This visualization provides an overview of the global distribution of research activities. More than the 70% of the authors are affiliated with institutions in China.

### v2

@@ Abstract
In recent years, spatio-temporal graph neural networks (GNNs) have attracted considerable interest in the field of time series analysis, due to their ability to capture, at once, dependencies among variables and across time points. The objective of this systematic literature review is hence to provide a comprehensive overview of the various modeling approaches and application domains of GNNs for time series classification and forecasting. A database search was conducted, and 281 journal papers were selected for a detailed examination of the current state-of-the-art in the field. This examination is intended to offer to the reader a comprehensive collection of proposed models, links to related source code, available datasets, benchmark models, and fitting results. All this information is hoped to assist researchers in future studies. To the best of our knowledge, this is the first systematic literature review presenting a detailed comparison of results of current spatio-temporal GNN models applied to different domains. In its final part, this review discusses current limitations and challenges in the application of spatio-temporal GNNs, such as comparability, reproducibility, explainability, poor information capacity, and scalability.
@@ 1 Introduction
In recent years, graph neural networks (GNNs) have emerged as a powerful class of artificial neural network models aimed at processing data that can be represented as graphs [Zheng2024]. GNNs are in fact particularly well-suited for a wide range of practical and engineering applications where data naturally lend themselves to be represented as graph structures, such as applications to transportation networks [review_traffico_jiang], image analysis [GNN_image], classification of molecular structure [GNN_molecule], and natural language processing [GNN_NLP]. The intuitive concept behind the GNN approach is that the nodes of a graph can be put in correspondence to objects or concepts, whereas the edges can be assumed to represent their mutual relationships [Scarselli2009]. Data for these quantities can then be processed at once by a GNN model. GNNs can be used at three different graph levels, related to three distinct classes of problems: at graph level, at edge level, and at node level [review_traffico_con_codici]. In graph-level problems, the goal is to predict a property of a graph based on its entire structure, rather than on individual nodes or edges. For example, a molecule from a sample can be represented as a graph, and global properties of this molecule can be inferred from data coming from its entire structure in relation to the sample. In edge-level problems, the goal is to predict the presence or absence of edges between pairs of nodes. As an example, this task can be used in recommendation systems to predict potential connections between users and items on the basis of past interactions. In node-level problems, the goal is to predict the role or value of each node within a graph, in order to solve either a classification or a regression task. Noticeably, regression tasks can sometimes be put in the form of temporal “autoregressions”, that is, they can include time.
@@
The inclusion of time in GNN modeling has proven to have far-reaching consequences. For this reason, this review examines the available journals’ literature on the specific use of GNNs on time series related tasks, with a particular focus on time series classification and forecasting. The idea behind the use of GNNs for time series problems is that GNNs can be made capable of capturing complex relationships, both inter-variable (connections between different variables within a multivariate series) but also inter-temporal (dependencies between different points in time) at once [review_Alippi]. The inclusion of this capability in GNN models results in a spatio-temporal GNN approach. Within this approach, the spatial dimension is related to the multivariate framework, where nodes typically correspond to different features of a multivariate time series at a given time step $t$, and edges to their relationships, and the temporal dimension is related to the temporal nature of the data. Namely, spatio-temporal GNNs are a class of GNN models designed to handle data in both spatial and temporal dimensions at once. In a spatio-temporal graph, data in each node or edge can evolve over time, and the challenge is then to model their related dynamic interactions over time.
@@
Notwithstanding all the potential interest for such important features, to the best of our knowledge there is a lack of systematic literature reviews (SLRs) on spatio-temporal GNN models for time series applications. Most of the existing surveys are not SLRs, or they are not recent, or only focus on a limited number of application domains and specific issues not necessarily related to time series. For example, recent reviews focus on algorithmic features [cit_rev2, cit_rev3, survey_Lucheroni, survey_Chikwendu], GNN characterization [cit_rev8, review_Alippi], specific fields [cit_rev4], and distributed training [cit_rev5]. Only one SLR on spatio-temporal GNNs was actually recently published, by [survey_spatiotemporal]. However, due to the restrictive search query used in that paper, that SLR covers 52 publications only. In addition, [survey_kumar_spatiotemporal_data] proposed a recent survey on spatio-temporal predictive modeling techniques for different domains. However, the authors approach the topic from a more general perspective, analyzing different machine learning models without placing too much emphasis on spatio-temporal GNNs. As an organized summary of these contributions, a list with a short description of the main reviews on spatio-temporal GNN models is provided in Tab. LABEL:t:tabella_altre_review.
Table 1: Short description of main reviews on spatio-temporal GNN models.
Review Description
Wu et al. (2020) [cit_rev3] One of the first reviews on the use of GNNs in machine learning and data mining. It introduces a taxonomy that classifies GNN architectures into four categories: recurrent, convolutional, graph autoencoders, and spatial-temporal GNNs. The focus of this review is primarily on early GNN models, rather than applications on time series data.
Zhou et al. (2020) [review_traffico_con_codici] The authors outline a general design pipeline for GNN models, covering key aspects such as the graph structure, pooling techniques, loss function design, and components of the propagation module (including recurrent and convolutional operators, and skip connections). They present some applications of GNNs, including physical systems modeling, chemical analysis, text processing, image classification, recommendation systems, and analysis of other types of data.
Jin et al. (2023) [cit_rev4] Review on the applications of spatio-temporal GNNs to predictive learning in urban computing. In the specific, the authors describe four methods for the construction of pre-defined graphs: topology-based, distance-based, similarity-based and interaction-based approaches. Then, they discuss various design strategies for spatio-temporal GNNs and their combination with some advanced learning frameworks, such as adversarial learning, continuous learning, and physics informed learning.
Li et al. (2023) [li2023graphneuralnetworkspatiotemporal] The authors aim to provide a systematic review of spatio-temporal mining using GNNs. After providing an overview of the graph structure and the spatial and temporal components of GNNs, the authors describe some applications in fields such as transportation, disaster management, environment, power systems, and neuroscience. However, unlike our review, they do not outline the selection process of the papers or include detailed and comprehensive tables that cover datasets, benchmarks and results.
Lin et al. (2023) [cit_rev5] This review is specialized on distributed training of GNNs. It investigates optimization techniques used in this context, including parallel mini-batch generation, dynamic mini-batch allocation, mini-batch transmission pipelining, and parallel aggregation with edge partitioning.
Longa et al. (2023) [cit_rev2] The authors offer an overview of GNNs for temporal graphs. They describe supervised and unsupervised learning tasks on temporal graphs, and present a taxonomy of temporal GNNs that groups the model in snapshot-based and event-based approaches.
Chen et al. (2024) [survey_Lucheroni] One of the first reviews on the use of GNNs for time series modeling. The authors thoroughly describe a wide range of models for time series classification, forecasting, and anomaly detection, based either on recurrent/convolutional GNNs and attentional GNNs. Moreover, they provide a brief comparison of the models’ performances in traffic forecasting and anomaly detection in water treatment.
Chikwendu et al. (2024) [survey_Chikwendu] This survey provides an overview of graph neural architecture search, and explores methods for representing nodes, edges, and sub-graphs, which add context and semantics to graphs. The authors also describe some applications, including computer vision, brain networks, recommendation system, biomedical application, natural language processing, and traffic.
Feng et al. (2024) [feng2024comprehensivesurveydynamicgraph] This survey examines 81 dynamic GNN models. It introduces a novel taxonomy that categorizes GNNs into discrete-time dynamic graphs and continuous-time dynamic graphs based on their structures, features, and dynamic modeling methods. In addition, the authors perform numerical experiments which evaluate 12 significant models on 6 graph datasets, which are not inherently related to time series.
Jin et al. (2024) [review_Alippi] The authors present a comprehensive review of GNNs for time series analysis, covering forecasting, classification, anomaly detection, and imputation. They categorize spatio-temporal GNNs according to their three main components: the spatial module, the temporal module, and the overall model architecture. Although they focus on the same topic as our SLR, they do not provide a detailed comparison of the results of existing models in the literature.
Khemani et al. (2024) [Khemani2024] The review discusses general concepts of GNNs, including architectures, techniques, datasets, applications, and challenges, offering a broad analysis that emphasizes these aspects without focusing specifically on time series analysis. It provides a brief description of papers in the literature, detailing application areas, datasets used, models applied, and performance evaluation.
Kumar et al. (2024) [survey_kumar_spatiotemporal_data] This SLR includes 200 papers from high-impact journals and prominent conferences on spatio-temporal predictive modeling techniques across different domains. Although its structure is similar to that of our review, its focus is broader and not specifically on spatio-temporal GNNs, as it also covers and discusses statistical and non-GNN machine learning models.
Wu et al. (2024) [cit_rev8] This review offers a comprehensive survey of GNN models from the perspective of computation theory. In particular, the authors compare existing GNN models through various perspectives: time, memory, communication, parallelism, and execution.
Zeghina et al. (2024) [survey_spatiotemporal] This SLR includes 52 papers on spatio-temporal GNNs across various application domains. Compared to our SLR, it covers a limited set of papers due to a restricted search query and it does not provide detailed and comprehensive tables of datasets, benchmarks, and model results.
@@
The intention of the SLR presented here is thus to provide a broad and comprehensive overview of the applications of spatio-temporal GNNs for time series classification and forecasting in different areas. Necessarily, this SLR also aims to assess if the current popularity of spatio-temporal GNN models is indeed due to their effectiveness, and to evaluate their efficacy and accuracy in different fields. Additionally, it collects results and benchmarks used in the discussed papers in order to assist researchers in their work by synthesizing the available literature on GNN applications. To the best of our knowledge, this is the first review presenting comprehensive tables with all the results from the various models and benchmarks proposed in a highly fragmented literature. This synthesis and comparison of the results is performed by gathering data from the existing literature, without conducting any numerical experiment on data. Moreover, given the large number of selected papers, it was not feasible to include a description of all the proposed models. To compensate this limitation, we provide a description of the most commonly used benchmarks and the most accurate models within each group. In case, for more detailed descriptions of experimental setups and data pre-processing techniques, the reader is referred to the original papers. It is hoped that this collected and distilled knowledge can anyway become a valuable resource and reference for researchers.
@@
During the development of this review, we formulated two sets of questions, one general and one specific, which are outlined below. The answers to the first set want to provide a general overview of the publications in the field, whereas the answers to the second set focus on specific aspects of the proposed GNN models.
@@ 1. GQ1)
    Trend: What is the temporal trend of the number of publications on GNN models? Is there a growing interest in GNNs?
@@ 3. GQ3)
    Journals: In which journals are the papers on GNNs published? Are these journals chosen because of the specific domain of application, or not?
@@ 3. SQ3)
    Taxonomy: Among the various taxonomy classes for GNNs, which are the most common? Are there some recurring mechanisms?
@@ 4. SQ4)
    Benchmark models: Which are the most commonly used benchmark models? Are they non-GNN machine learning models or also GNN models?
@@ 5. SQ5)
    Benchmark datasets: Which are the most commonly used benchmark datasets?
@@ 6. SQ6)
    Modeling paradigms: Which is the most common modeling paradigm? Is it modeling complex interacting systems, or modeling a system with multiple interacting quantities? Specifically, do graphs typically condense the relationships among multiple entities, or they describe different aspects of the same entity? In other words, is the used graph structure homogeneous or heterogeneous?
@@ 7. SQ7)
As the discussion will progress, the answers to both the general and the specific questions will be integrated in it. This will hopefully provide a comprehensive overview of the state-of-the-art of spatio-temporal GNN models. At the same time, a general overview of such GNN models will be assembled by highlighting similarities and differences between the approaches across different fields. In case of need of an even more detailed analysis of a specific discipline, or other aspects of GNN models, the reader is invited to consult the specialized reviews mentioned before [cit_rev1, cit_rev2, cit_rev3, survey_Lucheroni, cit_rev8, cit_rev4, cit_rev5, cit_rev6].
The paper is organized as follows. After this Introduction, Sec. 2 illustrates the methodology used to collect the papers. Sec. 3 provides a general overview of the selected papers. Sec. 4 introduces some fundamental definitions and notions at the basis of spatio-temporal GNN models, and discusses their taxonomy and the determination of the graph structure. Sec. 5 is the core of this review. It explores the diverse domains of application of the spatio-temporal GNN models proposed in the selected papers for the identified domains. A discussion of the findings and the answers to the research questions mentioned above are provided in Sec. 6. Finally, limits, challenges and future research directions are presented in Sec. 7. Sec. 8 concludes the review. A provides a list of all journal papers included in this SLR, together with the year of publication, group they belong to, case study, and nature of the task (e.g., classification or forecasting). All links mentioned in the papers are included in this review. For the reader’s convenience, links that no longer work at the time of submission are marked with a .
@@ 2 Methodology for the systematic literature review
In order to conduct this SLR, four databases of primary importance were consulted: Scopus, IEEE Xplore, Web of Science, and ACM. These databases were chosen for their comprehensive coverage of academic and scientific literature, ensuring a broad and inclusive search across multiple disciplines.
Prior to commencing the literature search, some exclusion criteria were defined. First, papers from the current year 2025 were excluded because it has not yet ended, hence the review will only cover the literature up to the end of year 2024. Second, only journal articles were considered, whereas both book chapters and conference papers were excluded. Such a selection is made for two reasons. Whereas it is true that new research directions often emerge from conference papers, which can contain outstanding results, most often the results presented in this form are considered preliminary even by the authors themselves, frequently because felt as lacking the stability required for standard journal publication. In addition, the exclusion of conference papers is made for readability reasons, specifically to ensure that a manageable number of articles can be discussed in full within a limited number of pages. Nevertheless, conference papers that provide useful benchmarks and which are frequently referenced as accurate benchmark models in other journal articles, are anyway referenced throughout the text.
An advanced search query to be submitted to the databases was designed to capture the breadth of the topic while maintaining specificity. Its general form is (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)  [^1] Specific search queries for the different databases: Scopus: TITLE-ABS-KEY((("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting)) AND PUBYEAR < 2025 AND (EXCLUDE(DOCTYPE,"cr") OR EXCLUDE (DOCTYPE,"cp")). IEEE Xplore: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). Filters applied: Journals, 2000 - 2024. Web of Science: (("graph neural network" OR gnn) AND "time series") AND (classification OR forecasting). Filters applied: Article, 2000 - 2024. ACM: [[All: "graph neural network"] OR [All: gnn]] AND [All: "time series"] AND [[All: classification] OR [All: forecasting]] AND [E-Publication Date: (01/01/2000 TO 12/31/2024)]. . The first two elements of the query aim to restrict the pool of GNN models to the time series domain, and the second part focuses the search on classification and forecasting applications.
Overall, the selection process of the papers followed the multi-stage approach described below:
@@
    Data extraction: relevant data and features were extracted from the included studies, and analyzed in order to better discuss the approach and present the most common datasets and models for each identified field.
A graphical representation of this selection process is provided in Fig. 1. A total of 879 records were identified, 355 from Scopus, 67 from IEEE Xplore, 194 from Web of Science, and 263 from ACM. After that, 253 duplicates were removed. Then, the titles and abstracts of the remaining 626 records were manually evaluated to select the pertinent papers. Among these, 333 were discarded because they were considered outside the scope of this review, and 12 were discarded because they were not in English. Hence, 281 records were included in the review.
Figure 1: PRISMA flowchart summarizing the identification and selection process, adapted from [PRISMA_statement]. The flowchart is structured vertically into three key phases: identification (collection of records from databases and removal of duplicates), screening (assessment for eligibility and exclusion of out-of-scope studies), and inclusion (inclusion of papers in the review). The boxes on the right indicate the number of studies excluded at each stage.
As for conference papers, a total of 1,765 records were identified across the four included databases. Assuming the same proportion of duplicates and articles to be excluded as those found for journal papers, about 600 conference papers would remain, making the total number of papers excessive. As previously stated, these conference papers have been excluded from this review. Next, Sec. 3 provides an overview of the selected journal publications.
@@ 3 Overview of the publications
This section provides a comprehensive overview of the 281 selected journal papers. After their selection, the papers were grouped according to their domain of application, as better explained in the thematic analysis of Sec. 5. In the following, some overall bibliographic information about the reviewed articles is presented.
Fig. 2 displays the number of publications over time across different groups. A cumulative positive trend can be observed over time, reflecting an increasing interest in spatio-temporal GNN models across various fields, since 2020. The four most common groups that have seen the greatest increase in publications are “Environment”, “Generic”, “Mobility”, and “Predictive monitoring”. For more details, see also the group numerosity reported in Tab. 2.
@@
Field / Year 2020 2021 2022 2023 2024 Total
Energy 1 - 6 6 20 33
Environment - 2 4 21 16 43
Finance - - 3 6 7 16
Generic - 1 4 19 20 44
Health - 1 2 5 7 15
Mobility 2 4 15 15 30 66
Other - 2 4 14 9 29
Predictive monitoring - 2 3 11 19 35
Total 3 12 41 97 128 281
The catalog of the publication sources indicates that, in the studies period, 145 different journals have published articles on GNN models for time series forecasting or classification. Tab. 3 lists the journals with more than 5 published papers.
Table 3: Journals with more than 5 published paper.
Journal Papers
IEEE Access 14
ACM Trans. Knowl. Discov. Data 10
Knowledge-Based Systems 9
IEEE Transactions on Instrumentation and Measurement 8
Applied Intelligence 8
IEEE Transactions on Knowledge and Data Engineering 7
Expert Systems with Applications 7
ACM Trans. Intell. Syst. Technol. 7
Sensors 6
IEEE Transactions on Intelligent Transportation Systems 6
IEEE Transactions on Industrial Informatics 6
Applied Energy 6
The diversity of the journals indicates a vast range of potential applications for spatio-temporal GNN models in different fields. This suggests that researchers often select journals based on the specific domain of application. This allows them to reach targeted audiences, and contribute within their respective fields. Such diversity also underlines the interdisciplinary nature of GNNs research, clearly spanning multiple domains. As for the publishers, there are around 40 different publishers, and the ones with more than 5 papers are listed in Tab. 4.
Table 4: Publishers with more than 5 published paper.
Journal Papers
Elsevier 82
Institute of Electrical and Electronics Engineers Inc. 42
Association for Computing Machinery 27
Multidisciplinary Digital Publishing Institute (MDPI) 27
IEEE 18
IEEE Computer Society 13
Springer 10
Springer Science and Business Media 10
John Wiley and Sons Inc 5
In order to examine the degree of collaboration among researchers, a co-authorship network was generated by using the VOSviewer software [VOSviewer]. Fig. 3 depicts authors as nodes of the network, and collaborations as connections, revealing some research clusters. Fig. 3 shows two of these clusters enlarged.
(a) Map of the complete collaboration network.
(b) Enlargement on two different clusters.
In Fig. 3, the color of the nodes is related to the average publication year of an author, so that the evolution of research groups over time can be evaluated. Two key observations can be made. First, the number of clusters suggests that there are many distinct research groups, indicating a high degree of fragmentation within the community. Second, the color of the nodes indicates that some groups were active only at the beginning and have subsequently disappeared, while others have recently emerged, as indicated by their yellow color. Some authors have had various collaborations over time, as also shown in Fig. 3.
Fig. 4 shows a pie chart of how corresponding authors are distributed among different countries. This visualization provides an overview of the global distribution of research activities. More than the 70% of the authors are affiliated with institutions in China.